# Optimal approximation of infinite-dimensional holomorphic functions II: recovery from i.i.d. pointwise samples

### v2

@@ 4.5 Upper bounds in probability
Let $\bm{\alpha},\bm{\beta}$ satisfy Assumption 4.1 with constant $\tau>0$ and $\varrho=\varrho_{\bm{\alpha},\bm{\beta}}$ be the tensor-product Jacobi probability measure (4.3). Let $\zeta>0$, $10<\epsilon<$ 1, $m\geq 3$ and consider sample points $\bm{y}_{1},\ldots,\bm{y}_{m}\sim_{\mathrm{i.i.d.}}\varrho$. Then there exists a reconstruction map $\mathcal{T}:(\mathcal{U}\times\mathcal{V})^{m}\rightarrow L^{2}_{\varrho}(\mathcal{U};\mathcal{V})$ depending on $\zeta$, $\epsilon$, $\bm{\alpha}$ and $\bm{\beta}$ only such that
$\mathbb{P}\Bigg{(}$ $\sup_{f\in\mathcal{H}(\bm{b})}{\|f-\mathcal{T}(\{(\bm{y}_{i},f(\bm{y}_{i}))\}^{m}_{i=1})\|}_{L^{2}_{\varrho}(\mathcal{U};\mathcal{V})}\leq{C(\bm{b},p,\zeta,{\tau})}\cdot\left(\frac{m}{{\log^{4+\zeta}(m)}+\log(\epsilon^{-1})}\right)^{\frac{1}{2}-\frac{1}{p}}$
@@
As with Theorem 4.4, the term $\log^{4+\zeta}(m)$ could be replaced by $\log^{4}(m)g(m)$ for any increasing function $g:\mathbb{N}\rightarrow(0,\infty)$ with $g(m)\rightarrow+\infty$ as $m\rightarrow+\infty$.
Theorems 4.5 and 4.6 are of independent interest, but they are also crucial in obtaining the ‘in expectation’ bounds, Theorems 4.3 and 4.4. Therefore, the remainder of the paper involves first establishing Theorems 4.5 and 4.6 and then using them to prove Theorems 4.3 and 4.4.
@@ 5 Jacobi polynomial expansions and best polynomial approximation
The proofs of Theorems 4.5 and 4.6 involve various best approximation error bounds for holomorphic functions using Jacobi polynomials. Therefore, in this section we first introduce Jacobi polynomial expansions, then provide the such bounds. Specifically, we consider the classical best $s$-term approximation and the so-called weighted best $k$-term approximation and best $s$-term approximation in anchored sets.
@@ 5.1 Jacobi polynomial expansions
are the coefficients of $f$, which are elements of $\mathcal{V}$. We write $\bm{c}=\bm{c}^{\bm{\alpha},\bm{\beta}}=(c^{\bm{\alpha},\bm{\beta}}_{\bm{\nu}})_{\bm{\nu}\in\mathcal{F}}$ for the sequence of coefficients. Finally, given an index set $S\subseteq\mathcal{F}$, we define the truncated expansion $f_{S}$ of $f$ as
@@ 5.2 $\ell^{p}$ spaces, norms and best approximation of sequences
Notice that ${\left\|\cdot\right\|}_{{p},\bm{w};\mathcal{V}}$ coincides with the unweighted norm ${\left\|\cdot\right\|}_{{p};\mathcal{V}}$ for any $\bm{w}$ when $p=2$, or for any $p$ when $\bm{w}=\bm{1}$. We define the weighted cardinality of an index set $S\subseteq{\Lambda}$ as $|{S}|_{\bm{w}}=\sum_{\bm{\nu}\in{S}}w^{2}_{\bm{\nu}}$ and, for $0<k\leq|\Lambda|_{\bm{w}}$, we define the $\ell^{p}_{\bm{w}}$-norm weighted best $(k,\bm{w})$-term approximation error of a $\mathcal{V}$-valued vector or sequence $\bm{c}\in\ell^{p}_{\bm{w}}(\Lambda;\mathcal{V})$ as
@@
We also need several concepts involving anchored sets. A set $\Lambda\subseteq\mathbb{N}^{d}_{0}$, where $d\in\mathbb{N}\cup\{\infty\}$, of (multi-)indices is lower if, whenever $\bm{\nu}\in\Lambda$ and $\bm{\mu}\leq\bm{\nu}$, it also holds that $\bm{\mu}\in\Lambda$. Moreover, $\Lambda$ is anchored if it is lower and if, whenever $\bm{e}_{j}\in\Lambda$ for some $j\in[d]$, it also holds that $\{\bm{e}_{1},\bm{e}_{2},\ldots,\bm{e}_{j}\}\subseteq\Lambda$.
A scalar-valued sequence $\bm{d}=(d_{\bm{\nu}})_{\bm{\nu}\in\Lambda}$ is monotonically nonincreasing if $d_{\bm{\nu}}\geq d_{\bm{\mu}}$ whenever $\bm{\nu}\leq\bm{\mu}$. It is anchored if it is monotonically nonincreasing and $d_{\bm{e}_{j}}\leq d_{\bm{e}_{i}}$ whenever $i,j\in[d]$ with $i\leq j$. Now let $\bm{c}\in\ell^{\infty}({\Lambda};\mathcal{V})$. An anchored majorant of $\bm{c}$ is any scalar-valued sequence $\bm{d}=(d_{\bm{\nu}})_{\bm{\nu}\in\Lambda}$ that is anchored and satisfies $d_{\bm{\nu}}\geq{\|c_{\bm{\nu}}\|}_{\mathcal{V}}$, $\forall\bm{\nu}\in\Lambda$. The minimal anchored majorant of $\bm{c}$ is the (unique) scalar-valued sequence $\tilde{\bm{c}}$ that is an anchored majorant and for which $\tilde{\bm{c}}\leq\bm{d}$ for any other anchored majorant $\bm{d}$. It has the explicit expression
@@ 5.3 Best approximation of polynomial coefficients
    There exists a set $S\subset\mathcal{F}$ with $|S|_{\bm{u}}\leq k$ depending on $\bm{b}$, $\bm{\alpha}$ and $\bm{\beta}$ only such that
@@
    There exists an anchored set $S\subset\mathcal{F}$ with $|S|\leq s$ depending on $\bm{b}$, $\bm{\alpha}$, $\bm{\beta}$ and $q$ only such that
@@
    where $\bm{c}=(c_{\bm{\nu}})_{\bm{\nu}\in\mathcal{F}}$ are the coefficients (5.2) of $f$ with respect to the tensor Jacobi polynomials (4.5). Moreover, the set can be chosen to be nested. If $s^{\prime}\geq s$, then there are sets $S\subseteq S^{\prime}\subset\mathcal{F}$ with $|S|\leq s$ and $|S^{\prime}|\leq s^{\prime}$ such that (5.10) holds for $S$ with $s$ and $S^{\prime}$ with $s^{\prime}$.
@@
These two theorems are extensions of well-known results in the literature. Early results of this type were developed in [11, 13, 14] for Taylor or Legendre polynomial coefficients of holomorphic functions arising as solution maps of certain parametric PDEs and considering, primarily, the best $s$-term approximation (5.4), the best $s$-term approximation in anchored sets (5.7) or a variant of the latter based on lower sets only. A more direct proof for Chebyshev or Legendre polynomial coefficients of $(\bm{b},\varepsilon)$-holomorphic functions (as opposed to parametric PDE solutions) was presented in [5, Thms. 3.28, 3.33], based on the arguments in [13]. Analysis of the best $(k,\bm{u})$-term approximation error was first considered in [5, Lem. 7.23] and later in [4, Thm. A.3], again for Chebyshev or Legendre polynomial coefficients only.
Parts (a) of Theorems 5.1 and 5.2 extend these previous works by allowing for arbitrary Jacobi polynomial coefficients. We also make explicit the fact that the sets can be chosen in a nested fashion, as this will be needed in later proofs. The uniform bounds for the constants given in parts (b) of Theorems 5.1 and 5.2 are novel, and were not considered in these previous works. They generalize results first shown in [6], which considered Legendre polynomials only.
@@ 5.4 Bounds for Jacobi polynomial coefficients of holomorphic functions and an abstract summability lemma
The remainder of this section is devoted to the proofs of Theorems 5.1 and 5.2. They involve two main ingredients. First, a bound for the Jacobi polynomial coefficients of $(\bm{b},\varepsilon)$-holomorphic functions, and second, an abstract summability lemma for sequences satisfying such a bound. Using these results, to prove Theorem 5.1 we construct a sequence $\bar{\bm{d}}$ based on the coefficients $\bm{c}$ and weights $\bm{u}$, and show that it is $\ell^{p}$-summable. For Theorem 5.2 we construct a similar sequence $\bar{\bm{d}}$ and show that it is $\ell^{p}_{\mathsf{A}}$-summable. Finally, the proofs of Theorems 5.1 and 5.2 follow by considering appropriate versions of Stechkin’s inequality.
We commence with the coefficient bound. Recall that $\mathcal{F}$ is the set of multi-indices with at most finitely-many nonzero entries as in (4.4). Consider $\bm{\nu}\in\mathcal{F}$. Then there exists $N\in\mathbb{N}$ such that $\nu_{k}=0$ for all $k>N$. Thus, the following lemma is a straightforward extension of [41, Cor. B.2.7]. For this reason, we omit its proof.
@@
Let $\bm{\alpha},\bm{\beta}>-\bm{1}$ and suppose that $f:\mathcal{U}\rightarrow\mathcal{V}$ is holomorphic in $\mathcal{E}_{\bm{\rho}}$ for some $\bm{\rho}\geq\bm{1}$. Then its coefficients $\bm{c}=(c_{\bm{\nu}})_{\bm{\nu}\in\mathcal{F}}$ in (5.2) with respect to the tensor Jacobi polynomials (4.5) satisfy
@@
Proof of (5.15).
Write $\mathbb{N}=E\cup F$, where $E=[d]$ and $F=\mathbb{N}\backslash[d]$ and $d\in\mathbb{N}$ is to be chosen later. Define $\kappa=1+{\varepsilon}/({2{\|\bm{b}\|}_{1}})$ and, for $\bm{\nu}\in\mathcal{F}$, the sequence $\tilde{\bm{\rho}}=\tilde{\bm{\rho}}(\bm{b},\varepsilon,\bm{\nu})\geq\bm{1}$ by
@@
Then it is readily checked that $\bm{\rho}=\tilde{\bm{\rho}}$ satisfies (5.14) and $\{k:\rho_{k}>1\}=\mathrm{supp}(\bm{\nu})$. Moreover, using the fact that the solution $x=a+\sqrt{a^{2}-1}$ to the equation $(x+1/x)/2=a$ satisfies $x\geq a$, we see that $\tilde{\rho_{j}}\geq\kappa>1$, $\forall j\in\mathrm{supp}(\bm{\nu})$. Since the function $\xi$ is continuous on $(1,\infty)$ and bounded as $t\rightarrow\infty$, there is a constant $C(\kappa,\xi)$ such that
@@
Proof of (5.17).
We write $\mathbb{N}=E\cup F$, where $E=[d]$ and $F=\mathbb{N}\backslash[d]$ once more, where $d\in\mathbb{N}$ is now large enough so that
@@
for some $\beta$ that will be chosen later. We now define $\tilde{\bm{\rho}}=\tilde{\bm{\rho}}(\bm{b},\beta,\varepsilon,\bm{\nu})\geq\bm{1}$ by
@@
We first construct $\widetilde{B}(\bm{\nu})$. We split $B(\bm{\nu})=B_{E}(\bm{\nu})B_{F}(\bm{\nu})$ once more, where $B_{S}(\bm{\nu})$ is as in (5.21) with this value of $\tilde{\bm{\rho}}$. Now consider $B_{E}(\bm{\nu})$. Note that there is a constant $D=D(\tilde{c},\tilde{\gamma},\tilde{\kappa})\geq 1$ such that
@@
$B_{F}(\bm{\nu})$ $\leq\prod_{j\in\mathrm{supp}(\bm{\nu}_{F})}{(\tilde{c}\nu_{j}+1)^{\tilde{\gamma}}}\left(\beta+\frac{\varepsilon\nu_{j}}{4b_{j}\|\bm{\nu}_{F}\|_{1}}\right)^{-\nu_{j}}\leq{\prod_{j\in\mathrm{supp}(\bm{\nu}_{F})}(\tilde{c}\nu_{j}+1)^{\tilde{\gamma}}\left(\beta+\frac{\varepsilon\nu_{j}}{4\tilde{b}_{j}\|\bm{\nu}_{F}\|_{1}}\right)^{-\nu_{j}}=:\widetilde{B}_{F}(\bm{\nu}).}$
Here $\tilde{\bm{b}}=(\tilde{b}_{j})_{j\in\mathbb{N}}\in\ell^{p}(\mathbb{N})$ is the minimal monotone majorant of $\bm{b}$, as defined in §2.2. Having specified both terms, we now formulate the upper bound $\widetilde{B}(\bm{\nu})=\widetilde{B}_{E}(\bm{\nu})\widetilde{B}_{F}(\bm{\nu})$.
We next show that $\widetilde{B}(\bm{\nu})$ is monotonically nonincreasing. Following Step 3 of the proof of [5, Thm. 3.33], we do this by showing that this holds for both $\widetilde{B}_{E}(\bm{\nu})$ and $\widetilde{B}_{F}(\bm{\nu})$ for $\bm{\nu}\in\mathcal{F}$. Let $\bm{\mu}\leq\bm{\nu}$ and note that $\mathrm{supp}(\bm{\bm{\mu}}_{E})\subseteq\mathrm{supp}(\bm{\nu}_{E})$. Then, since $\eta={(1+\tilde{\kappa})}/{(2\tilde{\kappa})}<1$, we readily see that $\widetilde{B}_{E}(\bm{\nu})\leq\widetilde{B}_{E}(\bm{\mu})$. To prove monotonicity of $\widetilde{B}_{F}(\bm{\nu})$ we show that for every $\bm{\nu}\in\mathcal{F}$ and every $i>d$ the inequality $\widetilde{B}_{F}(\bm{\nu}+\bm{e}_{i})\leq\widetilde{B}_{F}(\bm{\nu})$ holds. First, assume that $\nu_{i}\neq 0$. Using the definition of $\widetilde{B}_{F}(\bm{\nu})$, we see that
@@
$\dfrac{\widetilde{B}_{F}(\bm{\nu}+\bm{e}_{i})}{\widetilde{B}_{F}(\bm{\nu})}$ $=\frac{(\tilde{c}(\nu_{i}+1)+1)^{\tilde{\gamma}}}{(\tilde{c}\nu_{i}+1)^{\tilde{\gamma}}}\frac{\left(\beta+\frac{\varepsilon(\nu_{i}+1)}{4\tilde{b}_{i}({\|\bm{\nu}_{F}\|}_{1}+1)}\right)^{-(\nu_{i}+1)}}{\left(\beta+\frac{\varepsilon\nu_{i}}{4\tilde{b}_{i}{\|\bm{\nu}_{F}\|}_{1}}\right)^{-\nu_{i}}}\prod_{j\in\mathrm{supp}(\bm{\nu}_{F})\backslash\{i\}}\left(\frac{\beta+\frac{\varepsilon\nu_{j}}{4\tilde{b}_{j}({\|\bm{\nu}_{F}\|}_{1}+1)}}{\beta+\frac{\varepsilon\nu_{j}}{4\tilde{b}_{j}{\|\bm{\nu}_{F}\|}_{1}}}\right)^{-\nu_{j}}$

### v3

@@ 4.4 Upper bounds in probability
Let $\bm{\alpha},\bm{\beta}$ satisfy Assumption 4.1 with constant $\tau>0$ and $\varrho=\varrho_{\bm{\alpha},\bm{\beta}}$ be the tensor-product Jacobi probability measure (4.3). Let $10<\epsilon<$ 1, $m\geq 3$ and consider sample points $\bm{y}_{1},\ldots,\bm{y}_{m}\sim_{\mathrm{i.i.d.}}\varrho$. Then there exists a reconstruction map $\mathcal{T}:(\mathcal{U}\times\mathcal{V})^{m}\rightarrow L^{2}_{\varrho}(\mathcal{U};\mathcal{V})$ depending on $\epsilon$, $\bm{\alpha}$ and $\bm{\beta}$ only such that
$\mathbb{P}\Bigg{(}$ $\sup_{f\in\mathcal{H}(\bm{b})}{\|f-\mathcal{T}(\{(\bm{y}_{i},f(\bm{y}_{i}))\}^{m}_{i=1})\|}_{L^{2}_{\varrho}(\mathcal{U};\mathcal{V})}\leq C(\bm{b},p,\tau)\cdot\left(\frac{m}{\log^{5}(m)+\log(\epsilon^{-1})}\right)^{\frac{1}{2}-\frac{1}{p}}$
@@
As with Theorem 4.3, the term $\log^{5}(m)$ could be replaced by $\log^{4}(m)g(m)$ for any function $g$ satisfying the conditions detailed in Remark 4.3.
Theorems 4.5 and 4.6 are of independent interest, but they are also crucial in obtaining the ‘in expectation’ bounds, Theorems 4.2 and 4.3. Therefore, the remainder of the paper involves first establishing Theorems 4.5 and 4.6 and then using them to prove Theorems 4.2 and 4.3.
@@ 5 Jacobi polynomial expansions and best polynomial approximation
The proofs of Theorems 4.5 and 4.6 involve various best approximation error bounds for holomorphic functions using Jacobi polynomials. Therefore, in this section we first introduce Jacobi polynomials and Jacobi polynomial expansions, then provide such bounds. Specifically, we consider the classical best $s$-term approximation and the so-called weighted best $k$-term approximation and best $s$-term approximation in anchored sets.
5.1 Jacobi polynomials
  • Remark 5.1
    for some polynomial $p$ that is independent of $j$ and depends on $\bm{\alpha}$ and $\bm{\beta}$ only. See [7, Rmk. 3.5] for a similar assumption in a related context. In other words, Assumption 4.1 ensures that the $L^{\infty}$-norms of the constituent Jacobi polynomials do not blow up too fast. This is a key ingredient in our proofs, which involve showing various weighted best approximation error bounds for Jacobi polynomial expansions of $(\bm{b},1)$-holomorphic functions (see Theorems 5.2 and 5.3), where the weights are based on these norms (see (5.10)).
@@ 5.2 Jacobi polynomial expansions
are the coefficients of $f$, which are elements of $\mathcal{V}$. We write $\bm{c}=\bm{c}^{\bm{\alpha},\bm{\beta}}=(c^{\bm{\alpha},\bm{\beta}}_{\bm{\nu}})_{\bm{\nu}\in\mathcal{F}}$ for the sequence of coefficients. Next, given an index set $S\subseteq\mathcal{F}$, we define the truncated expansion $f_{S}$ of $f$ as
@@ 5.3 $\ell^{p}$ spaces, norms and best approximation of sequences
Notice that ${\left\|\cdot\right\|}_{p,\bm{w};\mathcal{V}}$ coincides with the unweighted norm ${\left\|\cdot\right\|}_{p;\mathcal{V}}$ for any $\bm{w}$ when $p=2$, or for any $p$ when $\bm{w}=\bm{1}$. Next, we define the weighted cardinality of an index set $S\subseteq{\Lambda}$ as $|S|_{\bm{w}}=\sum_{\bm{\nu}\in S}w^{2}_{\bm{\nu}}$ and, for $0<k\leq|\Lambda|_{\bm{w}}$, we define the $\ell^{p}_{\bm{w}}$-norm weighted best $(k,\bm{w})$-term approximation error of a $\mathcal{V}$-valued vector or sequence $\bm{c}\in\ell^{p}_{\bm{w}}(\Lambda;\mathcal{V})$ as
@@
We also need several concepts involving anchored sets. A set $\Lambda\subseteq\mathbb{N}^{d}_{0}$, where $d\in\mathbb{N}\cup\{\infty\}$, of (multi-)indices is lower if, whenever $\bm{\nu}\in\Lambda$ and $\bm{\mu}\leq\bm{\nu}$, it also holds that $\bm{\mu}\in\Lambda$. Moreover, $\Lambda$ is anchored if it is lower and if, whenever $\bm{e}_{j}\in\Lambda$ for some $j\in[d]$, it also holds that $\{\bm{e}_{1},\bm{e}_{2},\ldots,\bm{e}_{j}\}\subseteq\Lambda$. A scalar-valued sequence $\bm{d}=(d_{\bm{\nu}})_{\bm{\nu}\in\Lambda}$ is monotonically nonincreasing if $d_{\bm{\nu}}\geq d_{\bm{\mu}}$ whenever $\bm{\nu}\leq\bm{\mu}$. It is anchored if it is monotonically nonincreasing and $d_{\bm{e}_{j}}\leq d_{\bm{e}_{i}}$ whenever $i,j\in[d]$ with $i\leq j$. Now let $\bm{c}\in\ell^{\infty}({\Lambda};\mathcal{V})$. An anchored majorant of $\bm{c}$ is any scalar-valued sequence $\bm{d}=(d_{\bm{\nu}})_{\bm{\nu}\in\Lambda}$ that is anchored and satisfies $d_{\bm{\nu}}\geq{\|c_{\bm{\nu}}\|}_{\mathcal{V}}$, $\forall\bm{\nu}\in\Lambda$. The minimal anchored majorant of $\bm{c}$ is the (unique) scalar-valued sequence $\tilde{\bm{c}}$ that is an anchored majorant and for which $\tilde{\bm{c}}\leq\bm{d}$ for any other anchored majorant $\bm{d}$. It has the explicit expression
@@ 5.4 Best approximation of polynomial coefficients
    There exists a set $S\subset\mathcal{F}$ with $|S|_{\bm{u}}\leq k$ depending on $k$, $\bm{b}$, $\bm{\alpha}$ and $\bm{\beta}$ only such that
@@
    There exists an anchored set $S\subset\mathcal{F}$ with $|S|\leq s$ depending on $s$, $\bm{b}$, $\bm{\alpha}$, $\bm{\beta}$ and $q$ only such that
@@
    where $\bm{c}=(c_{\bm{\nu}})_{\bm{\nu}\in\mathcal{F}}$ are the coefficients (5.4) of $f$ with respect to the tensor Jacobi polynomials (5.2). Moreover, the sets can be chosen to be nested. If $s^{\prime}\in\mathbb{N}$, $s^{\prime}\geq s$, then there are sets $S\subseteq S^{\prime}\subset\mathcal{F}$ with $|S|\leq s$ and $|S^{\prime}|\leq s^{\prime}$ such that (5.12) holds for $S$ with $s$ and $S^{\prime}$ with $s^{\prime}$.
@@
These two theorems are extensions of well-known results in the literature. Early results of this type were developed in [11, 13, 14] for Taylor or Legendre polynomial coefficients of holomorphic functions arising as solution maps of certain parametric PDEs and considering, primarily, the best $s$-term approximation (5.6), the best $s$-term approximation in anchored sets (5.9) or a variant of the latter based on lower sets only. A more direct proof for Chebyshev or Legendre polynomial coefficients of $(\bm{b},\varepsilon)$-holomorphic functions (as opposed to parametric PDE solutions) was presented in [5, Thms. 3.28, 3.33], based on the arguments in [13]. Analysis of the best $(k,\bm{u})$-term approximation error was first considered in [5, Lem. 7.23] and later in [3, Thm. A.3], again for Chebyshev or Legendre polynomial coefficients only.
Parts (a) of Theorems 5.2 and 5.3 extend these previous works by allowing for arbitrary Jacobi polynomial coefficients. We also make explicit the fact that the sets can be chosen in a nested fashion, as this will be needed in later proofs. The uniform bounds for the constants given in parts (b) of Theorems 5.2 and 5.3 are novel, and were not considered in these previous works. They generalize results first shown in [6], which considered Legendre polynomials only.
@@ 5.5 Bounds for Jacobi polynomial coefficients of holomorphic functions and an abstract summability lemma
The remainder of this section is devoted to the proofs of Theorems 5.2 and 5.3. They involve two main ingredients. First, a bound for the Jacobi polynomial coefficients of holomorphic functions, and second, an abstract summability lemma for sequences satisfying such a bound. Using these results, to prove Theorem 5.2 we construct a sequence $\bar{\bm{d}}$ based on the coefficients $\bm{c}$ and weights $\bm{u}$, and show that it is $\ell^{p}$-summable. For Theorem 5.3 we construct a similar sequence $\bar{\bm{d}}$ and show that it is $\ell^{p}_{\mathsf{A}}$-summable. Finally, the proofs of Theorems 5.2 and 5.3 follow by considering appropriate versions of Stechkin’s inequality.
Consider $\bm{\nu}\in\mathcal{F}$, where $\mathcal{F}$ is as in (5.1). Then there exists $N\in\mathbb{N}$ such that $\nu_{k}=0$ for all $k>N$. Thus, the following lemma is a straightforward extension of [41, Cor. B.2.7]. For this reason, we omit its proof.
@@
Let $\bm{\alpha},\bm{\beta}>-\bm{1}$ and suppose that $f:\mathcal{U}\rightarrow\mathcal{V}$ is holomorphic in $\mathcal{E}_{\bm{\rho}}$ for some $\bm{\rho}\geq\bm{1}$. Then its coefficients (5.4) with respect to the tensor Jacobi polynomials (5.2) satisfy
@@
Proof of (5.17). Write $\mathbb{N}=E\cup F$, where $E=[d]$ and $F=\mathbb{N}\backslash[d]$ and $d\in\mathbb{N}$ is to be chosen later. Define $\kappa=1+{\varepsilon}/({2{\|\bm{b}\|}_{1}})$ and, for $\bm{\nu}\in\mathcal{F}$, the sequence $\tilde{\bm{\rho}}=\tilde{\bm{\rho}}(\bm{b},\varepsilon,\bm{\nu})\geq\bm{1}$ by
@@
Then it is readily checked that $\bm{\rho}=\tilde{\bm{\rho}}$ satisfies (5.16) and $\{k:\rho_{k}>1\}=\mathrm{supp}(\bm{\nu})$. Moreover, using the fact that the solution $x=a+\sqrt{a^{2}-1}$ to the equation $(x+1/x)/2=a$, where $a\geq 1$, satisfies $x\geq a$, we see that $\tilde{\rho_{j}}\geq\kappa>1$, $\forall j\in\mathrm{supp}(\bm{\nu})$. Since the function $\xi$ is continuous on $(1,\infty)$ and bounded as $t\rightarrow\infty$, there is a constant $C(\kappa,\xi)$ such that
@@
Proof of (5.19). We write $\mathbb{N}=E\cup F$, where $E=[d]$ and $F=\mathbb{N}\backslash[d]$ once more and $d\in\mathbb{N}$ is now large enough so that
@@
for some $\beta$ that will be chosen later. Given $\bm{\nu}\in\mathcal{F}$, we now define $\tilde{\bm{\rho}}=\tilde{\bm{\rho}}(\bm{b},\beta,\varepsilon,\bm{\nu})\geq\bm{1}$ by
@@
We first construct $\widetilde{B}(\bm{\nu})$. We split $B(\bm{\nu})=B_{E}(\bm{\nu})B_{F}(\bm{\nu})$ once more, where $B_{S}(\bm{\nu})$ is as in (5.23) with $\tilde{\bm{\rho}}$ as defined by (5.30). Now consider $B_{E}(\bm{\nu})$. Note that there is a constant $D=D(\tilde{c},\tilde{\gamma},\tilde{\kappa})\geq 1$ such that
@@
$B_{F}(\bm{\nu})$ $\leq\prod_{j\in\mathrm{supp}(\bm{\nu}_{F})}(\tilde{c}\nu_{j}+1)^{\tilde{\gamma}}\left(\beta+\frac{\varepsilon\nu_{j}}{4b_{j}\|\bm{\nu}_{F}\|_{1}}\right)^{-\nu_{j}}$
@@
$\leq\prod_{j\in\mathrm{supp}(\bm{\nu}_{F})}(\tilde{c}\nu_{j}+1)^{\tilde{\gamma}}\left(\beta+\frac{\varepsilon\nu_{j}}{4\tilde{b}_{j}\|\bm{\nu}_{F}\|_{1}}\right)^{-\nu_{j}}=:\widetilde{B}_{F}(\bm{\nu}).$
Here $\tilde{\bm{b}}=(\tilde{b}_{j})_{j\in\mathbb{N}}\in\ell^{p}(\mathbb{N})$ is the minimal monotone majorant of $\bm{b}$, as defined in §2.2. Having specified both terms $\widetilde{B}_{E}(\bm{\nu})$ and $\widetilde{B}_{F}(\bm{\nu})$, we now formulate the upper bound $\widetilde{B}(\bm{\nu})=\widetilde{B}_{E}(\bm{\nu})\widetilde{B}_{F}(\bm{\nu})$.
We next show that $\widetilde{B}(\bm{\nu})$ is monotonically nonincreasing. Following Step 3 of the proof of [5, Thm. 3.33], we do this by showing that this holds for both $\widetilde{B}_{E}(\bm{\nu})$ and $\widetilde{B}_{F}(\bm{\nu})$. First, let $\bm{\mu}\leq\bm{\nu}$ and note that $\mathrm{supp}(\bm{\bm{\mu}}_{E})\subseteq\mathrm{supp}(\bm{\nu}_{E})$. Then, since $\eta={(1+\tilde{\kappa})}/{(2\tilde{\kappa})}<1$, we readily see that $\widetilde{B}_{E}(\bm{\nu})\leq\widetilde{B}_{E}(\bm{\mu})$. Hence $\widetilde{B}_{E}(\bm{\nu})$ is monotonically nonincreasing. To establish the same for $\widetilde{B}_{F}(\bm{\nu})$ we show that, for every $\bm{\nu}\in\mathcal{F}$ and every $i>d$ the inequality $\widetilde{B}_{F}(\bm{\nu}+\bm{e}_{i})\leq\widetilde{B}_{F}(\bm{\nu})$ holds. First, assume that $\nu_{i}\neq 0$. Using the definition of $\widetilde{B}_{F}(\bm{\nu})$, we see that